OpenStudy (anonymous):

Find the first derivative of each of the following function and leave your answer in its simplest form: Q1. y=2x^-5 + 3(sqrt x) - 8/7 dy/dx = -10x^-6 + 3/2 x ^(-1/2) ?? ------------------------------------- Q2)). r=((t³-2)/(4t⁵+19)) U=(t^3)-2 v= 4t^5 +19 du/dx=3t^2 dv/dx=20t^4 (v(du/dx)- u(dv/dx))/v^2 (4t^5 +19)(3t^2)-((t^3)-2 )(20t^4) ------------------------------ ( 4t^5 +19)^2 12t^7+57t^2 - 20t^7 + 40t^4 ---------------------------- ( 4t^5 +19)^2 -8t^7+40t^4-57t^2 ------------------ ( 4t^5 +19)^2 t^2(-8t^5+40t^2-57) ------------------- ( 4t^5 +19)^2

OpenStudy (anonymous):

is my answers correct?

OpenStudy (anonymous):

@hartnn @UnkleRhaukus @jennychan12 @AccessDenied

OpenStudy (unklerhaukus):

the first one is right!

OpenStudy (anonymous):

how about Q2?

OpenStudy (unklerhaukus):

i can't see any mistake,

OpenStudy (anonymous):

was that the reason i been ignored for so many hours ><

OpenStudy (unklerhaukus):

maybe because your question isn't in \(\LaTeX\)

OpenStudy (anonymous):

anyway thank you uncle

OpenStudy (unklerhaukus):

i think you can simply the second one a bit

OpenStudy (unklerhaukus):

Q2)\[r=\frac{t^3-2}{4t^5+19}\] \[u=t^3-2\qquad v= 4t^5 +19\]\[\frac{du}{dx}=3t^2\qquad\frac{dv}{dx}=20t^4\] \[\begin{align*} \frac{dr}{dx}&=\frac{v(du/dx)- u(dv/dx)}{v^2}\\ &=\frac{(4t^5 +19)(3t^2)-(t^3-2 )(20t^4)}{( 4t^5 +19)^2}\\ &=\frac{12t^7+57t^2 - 20t^7 + 40t^4}{( 4t^5 +19)^2}\\ &=\frac{-8t^7+40t^4-57t^2}{( 4t^5 +19)^2}\\ &=\frac{t^2(-8t^5+40t^2-57)}{( 4t^5 +19)^2}\end{align*}\]

OpenStudy (anonymous):

ya!! that why my final answer is t^2(-8t^5+40t^2-57) ------------------- ( 4t^5 +19)^2

OpenStudy (unklerhaukus):

actually you won't be able to simplify, but now i can see one tiny minus sign error

OpenStudy (unklerhaukus):

can you spot it?

OpenStudy (unklerhaukus):

on the 57

OpenStudy (anonymous):

u refering to +57?

OpenStudy (unklerhaukus):

yeah